OpenStudy (96mertcank96):

Solve the System : 1/x - 3/y = -3 2/x - 1/y = 4

OpenStudy (academicgurusinc):

So isolate x for 1/x-3/y = -3 and you'll get x=-y/(3y-3) Next, substitute x=-y/(3y-3) into your equation to get your "y" value. So you'll get y=1/2 Finally, plug in y=1/2 into your original equation to your "x" and you'll get x=1/3 and y=1/2 as your final answer If you found this helpful, I encourage to subscribe to our youtube channel https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)

OpenStudy (96mertcank96):

Thank you so much for your help

OpenStudy (academicgurusinc):

Not a problem :)

OpenStudy (96mertcank96):

@AcademicGurusInc

OpenStudy (academicgurusinc):

Tag me whenever you need help. and please subscribe to our channel :)

OpenStudy (96mertcank96):

i will as soon as i am done studying i have a question about the second part of your answer i do not understand what i have to do

OpenStudy (academicgurusinc):

So this basic method is called substitution. Here's how it works: you pretty much find a variable and you take it all by itself. For example if you have x+2=4y... you'll take "x" all by itself so x=4y-2 .... then when you separated one variable from the entire equation you'll plug it back to the formula to find your other variable.

OpenStudy (96mertcank96):

wow that is really confusing thank you for your help though

OpenStudy (anonymous):

You still need help on this?

OpenStudy (96mertcank96):

yes please

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):